OpenStudy (anonymous):

Find the value(s) of c guaranteed by the Mean Value Theorem for Integrals for the function over the given interval. (Round your answer to four decimal places. f(x) = 8 x , [4, 9]

OpenStudy (anonymous):

Can someone please help me...

OpenStudy (anonymous):

There's this formula, it was 1/b-a, in your case, its 1/(9-4) = 1/5. Multiply it by the integral from 4-9.

OpenStudy (zarkon):

solve \[\frac{1}{b-a}\int\limits_{a}^{b}f(x)dx=f(c)\] for c

OpenStudy (anonymous):

lot of these today

OpenStudy (zarkon):

I haven't been on much...guess I missed all the fun ;)